# via Dodecaneso 35, 16146, Genova, Italy

### v1

@@ 5.1 The form of $G$ and $G_{pub}$
Since the map $T(x,y)=x+{\mathfrak{q}}(y)$ and its inverse $T_{y}^{-1}(x)=x-{\mathfrak{q}}(y)$ have degree $2$, by Remark 3.2 the degree of the $t$-twisted map $G$ is at most $4$. We analyze the monomials appearing in $G$ (and in $G_{pub}$) more closely.
@@
First, we notice that the terms appearing in $T(x,y)$ are of the form $x_{i}$, $y_{i}$, $y_{i}y_{j}$. Clearly, the same terms will appear in $T_{y}^{-1}(x)$, and so the degrees of the first $t$ coordinates of $F$ are the same of the degrees of the first $t$ coordinates of $G$. Now, we consider the last $m-t$ coordinates of $G$, corresponding to $U(T_{y}^{-1}(x),y)=U(x-{\mathfrak{q}}(y),y)$. Since $U$ consists of OV quadratic polynomials $f^{(i)}$ of the form of Equation (2), in $U(x,y)$ we can find terms of the form $x_{i},y_{i},x_{i}x_{j},x_{i}y_{j},y_{i}y_{j}$. By evaluating $x\mapsto x-{\mathfrak{q}}(y)$, the variable $x_{i}$ might produce terms of the form $x_{j}$, $y_{j}$ and $y_{j}y_{k}$. Therefore, potentially, in the last $m-t$ coordinates of $G$ we can have terms of the form $x_{i}$, $y_{i}$, $x_{i}x_{j}$, $x_{i}y_{j}$, $y_{i}y_{j}$, $x_{i}y_{j}y_{k}$, $y_{i}y_{j}y_{k}$, $y_{i}y_{j}y_{k}y_{l}$. Notice that, even if we replace $U$ with a dense quadratic map, the possible terms in $G$ have the same form. To sum up, even if the degree of $G$ is up to 4, the terms of degree 4 involve only variables in $y$, while terms of degree 3 involve only variables in $y$ or 2 variables in $y$ and 1 variable in $x$.
The public map $G_{pub}=A_{1}\circ G\circ A_{2}$ consists of dense polynomials of degree up to 4, since the random map $A_{2}$ will remove the above-mentioned restrictions. However, the affine transformation keeps invariant the amount of components of a fixed degree, so to analyse the possible degrees of the components of $G_{pub}$, we can directly consider the map $G(x,y)=(x-{\mathfrak{q}}(y),U(x-{\mathfrak{q}}(y),y))$. So, $G_{pub}$ has at least $q^{t}-1$ quadratic components, since the components $\lambda\cdot G$ with $\lambda=(v,0_{m-t})$ with nonzero $v\in{\mathbb{F}}_{q}^{t}$ are quadratic.
We present in Table 1 some computational experiments on the possible degree of the system $U(x-{\mathfrak{q}}(y),y)$. We observed that all the constructed systems $U(x-{\mathfrak{q}}(y),y)$ do not have quadratic components.
$q$ $n$ $m$ $t$ $s$#\## of systems $[\mathtt{d}_{3},\mathtt{d}_{4}]$
5 5 4 2 1
4
46
$[4,20]$
$[0,24]$
5 6 5 2 2
16
34
$[3,120]$
$[0,124]$
5 10 8 3 2
1
49
$[4,3120]$
$[0,3124]$
$2^{6}$ 5 4 2 1 50 $[0,4095]$
$2^{6}$ 6 5 2 2
1
49
$[36,262080]$
$[0,262143]$
Table 1: For each choice of parameters, number of components $\mathtt{d}_{3}$ of degree $3$ and $\mathtt{d}_{4}$ of degree $4$ for 50 randomly generates systems $U(x-{\mathfrak{q}}(y),y)$
@@ 5.2 Dimensions of the keys
In the following, we report an analysis of the sizes of the public and private keys.
@@
Recall that a polynomial of degree $r$ in $n$ variables with coefficients over ${\mathbb{F}}_{q}$ has $M_{n}(r)={n+r\choose r}$ terms. Notice that when $q\leq r$, we can use the field equations $x_{i}^{q}-x_{i}$ to lower the degrees of the polynomials over ${\mathbb{F}}_{q}$ and store less coefficients. $G_{pub}$ is a system of $m$ equations of degree at most 4 in $n$ variables, where each equation has $M_{n}(4)={n+4\choose 4}$ possible terms. Hence, for $G_{pub}$ we need to specify $m\cdot M_{n}(4)$ coefficients over ${\mathbb{F}}_{q}$.
@@
  • $A_{1}$ and $A_{2}$ are an affine bijections over ${\mathbb{F}}_{q}^{m}$ and ${\mathbb{F}}_{q}^{n}$ respectively: They correspond to invertible $m\times m$ and $n\times n$ matrices over ${\mathbb{F}}_{q}$ plus a constant vector. So this means $m^{2}+m+n^{2}+n$ coefficients over ${\mathbb{F}}_{q}$.
@@
To give a more precise idea to the reader, in Table 2 we report the amounts for public and secret keys for some small parameters. The sizes are in terms of the amount of coefficients in ${\mathbb{F}}_{q}$ to store.
$n$ $m$ $t$ $s$ amount for $pk$ amount for $sk$
5 4 2 1 504 106
6 5 2 2 1050 177
10 8 3 2 8008 545
Table 2: Amount of coefficients of $\mathbb{F}_{q}$ to store
@@ 5.3 Computational cost of the procedure
In this subsection, we present a tentative analysis on the computational cost of the proposed multivariate scheme. Notice that the application of this procedure reduces to the evaluation of polynomials of degree at most four and computing the solution of linear systems. Considering the direct evaluation of a polynomial, we have the following estimates.
@@
This is clearly an upper bound, since more efficient techniques might be used, see for example [BES13]. We indicate with $\mathfrak{M}(r,n)$ the number of multiplications needed to solve a linear system of $r$ equations in $n$ variables. Recall that $\mathfrak{M}(r,n)$ is of the order $rn\cdot\min(r,n)$, see for example [BV18, Appendix B].
@@
Consider a signature scheme based on the Pesto primitive as in Definition 4.2. Set $\mathfrak{m}_{i}(r)$ and $\mathfrak{M}(r,k)$ as defined before. Excluding the computation of the hash value, the computational cost, in terms of multiplications over ${\mathbb{F}}_{q}$, to verify the validity of a signature is $m\cdot\mathfrak{m}_{4}(n)$, whereas the computational cost to produce a valid signature is $m\cdot\mathfrak{m}_{1}(m)+t\cdot\mathfrak{m}_{2}(n-t)+n\cdot\mathfrak{m}_{1}(n)+(m-t)\left(\mathfrak{m}_{2}(t)+\mathfrak{m}_{2}(s)\right)+\mathfrak{M}(m-t,n-t-s)$.
@@
We are left with analysing the cost for the third step, where we have to find a preimage of $U(w_{T}^{\prime},y)=w_{U}^{\prime}$. Hence we need to set $y_{1},\ldots,y_{s}$ to random values, evaluate those variables and then solve a linear system. The cost of evaluating $s$ variables corresponds to $(m-t)\cdot\mathfrak{m}_{2}(s)=(m-t)s(s+2)$ multiplications. The cost of solving the final system of $m-t$ equations in $n-t-s$ variables is $\mathfrak{M}(m-t,n-t-s)$ multiplications.
@@ 6 Security Analysis
In this section, we provide some considerations on the security of the scheme Pesto of Definition 4.2. Among them, we consider attacks that have been exploited for the MI cryptosystem (and its generalizations) and we analyze under which conditions it would be possible to extend these also to our system.
@@ 6.1 The importance of $A_{2}$
We start this analysis by presenting a first observation on the security of the scheme, which stresses the importance of the affine bijection $A_{2}$. Indeed, if an attacker is able to recover $A_{2}$, then they can operate as follows. First, compute $\bar{G}=G_{pub}\circ A_{2}^{-1}$ and isolate its quadratic components: $t$ of them are of the form $x_{i}-{\mathfrak{q}}_{i}(y)$, for $i=1,\ldots,t$. From this, it is rather easy to solve the system. Suppose a cyphertext $c$ is given, and the attacker wants to find a message $m$ such that $G_{pub}(m)=c$. Write $G_{pub}(m)=\bar{G}\circ A_{2}(m)=c$. From the isolated quadratic equations, recover $x={\mathfrak{q}}(y)+c_{T}$. By substituting this into the remaining equations of $\bar{G}$, solve the system, which has at most $s$ variables appearing in quadratic terms. Notice that this last step is equivalent to what a legitimate user has to do. Once the solution $(x,y)$ of this system is found, then $m=A_{2}^{-1}(x,y)$.
In order to perform this attack, the attacker has to completely recover $A_{2}$ but only partially $A_{1}$ (it is enough to isolate the quadratic components of the form $x_{i}-{\mathfrak{q}}_{i}(y)$). The cost of isolate these quadratic components will be treated in the following section.
6.2 A partial recover of $A_{1}$ isolating the quadratic components
The goal of this attack is to isolate in $G_{pub}$ the components which correspond to (a linear combination of) the first $t$ coordinates of $G$. Notice that these components form an OV system which can be then attacked using other known techniques, later analyzed in §6.5.
We try to estimate the cost of isolating the quadratic components of $G_{pub}$. Let $r\geq t$ be such that the number of quadratic components of $G_{pub}$ is $q^{r}-1$. Notice that when $r=t$, then $U(x-{\mathfrak{q}}(y),y)$ does not have quadratic components. We consider the list $\Delta$ of all possible terms of degree 3 and degree 4 in $n$ variables appearing in $G_{pub}$. We have
@@ 6.2 A partial recover of $A_{1}$ isolating the quadratic components
We construct the matrix $A$ with $n$ rows and $|\Delta|$ columns, with at position $(i,j)$ the coefficients of the $j$-th term of $\Delta$ in the $i$-th equation of $G_{pub}$. So, to recover the quadratic equations of $G_{pub}$ it is enough to solve the linear system $x^{T}A=A^{T}x=0_{|\Delta|}$. The set of solutions corresponds to the set of quadratic components of $G_{pub}$. The size of this set is $q^{r}$. Recall that the cost of solving a linear system with $n$ equations in $|\Delta|$ variables $\mathfrak{M}(n,|\Delta|)$ is of the order $O(n^{6})$. Let us note that in this way we can recover the space of the quadratic components of $G_{pub}$, but not the exact quadratic coordinates of $G_{pub}$.
6.3 Differential attack: a partial recover of $A_{2}$ using the linear structures
@@
    $G_{pub}$ has (at least) $q^{t}-1$ quadratic components with (at least) $q^{t}$ linear structures in common (see Definition 6.1 below).
@@
    If we are able to identify this subspace, we know that the linear part of $A_{2}$ maps $V$ into ${\mathbb{F}}_{q}^{t}$. So, we can partially recover this linear part.
@@
The multiset of the number of linear structures of the components of $G$ and of $G_{pub}$ is the same. Moreover, $a\in{\mathbb{F}}_{q}^{n}$ is a linear structure of $\lambda\cdot G_{pub}$ if and only if $L_{2}(a)$ is a linear structure of ${L_{1}^{T}(\lambda)}\cdot G$ where $L_{1}$ and $L_{2}$ are the linear parts of the affine bijections $A_{1}$ and $A_{2}$, and $L_{1}^{T}$ is the linear function corresponding to the transpose of the matrix defining $L_{1}$, that is $x\cdot L_{1}(y)=L_{1}^{T}(x)\cdot y$ for every $x,y$.
Proof 6.5.
@@
The second statement can be deduced from the following. From the definition of $G_{pub}$ and since $A_{1}$ is affine we have $D_{a}(\lambda\cdot G_{pub})(x)=\lambda\cdot G_{pub}(x+a)-\lambda\cdot G_{pub}(x)=\lambda\cdot A_{1}\circ G\circ A_{2}(x+a)-\lambda\cdot A_{1}\circ G\circ A_{2}(x)=\lambda\cdot L_{1}\circ G\circ A_{2}(x+a)-\lambda\cdot L_{1}\circ G\circ A_{2}(x)$. Then $D_{a}(\lambda\cdot G_{pub})(x)=\lambda\cdot L_{1}(G(A_{2}(x)+L_{2}(a))-G(A_{2}(x)))=\lambda\cdot L_{1}(D_{L_{2}(a)}G(A_{2}(x)))=L_{1}^{T}(\lambda)\cdot D_{L_{2}(a)}G(A_{2}(x))$.
@@
The desired components of $G_{pub}$ are quadratic; therefore by isolating all the quadratic components of the public function, we are isolating the desired components plus (eventually) the quadratic components of $U(x-{\mathfrak{q}}(y),y)$. Notice that in all the computational experiments reported in Table 1 $U(x-{\mathfrak{q}}(y),y)$ did not have components of degree 2.

### v2

@@ 5.1 The form of $G$ and $G_{pub}$
Since the map $T(x,y)=x+{\mathfrak{q}}(y)$ and its inverse $T_{y}^{-1}(x)=x-{\mathfrak{q}}(y)$ have degree $2$, by Theorem 3.2 the degree of the $t$-twisted map $G$ is at most $4$. We analyse the monomials appearing in $G$ (and in $G_{pub}$) more closely.
@@
As stated above, the first $t$ coordinates of $G$ are those from $T_{y}^{-1}(x)=x-{\mathfrak{q}}(y)$, which have degree 2 and consist of terms of the form $x_{i},y_{i},y_{i}y_{j}$. Now, we consider the last $m-t$ coordinates of $G$, corresponding to $U(T_{y}^{-1}(x),y)=U(x-{\mathfrak{q}}(y),y)$. Since $U$ consists of OV quadratic polynomials $f^{(i)}$ of the form of Equation (4), in $U(x,y)$ we can find terms of the form $x_{i},y_{i},x_{i}x_{j},x_{i}y_{j},y_{i}y_{j}$. By evaluating $x\mapsto x-{\mathfrak{q}}(y)$, the variable $x_{i}$ might produce terms of the form $x_{j}$, $y_{j}$ and $y_{j}y_{k}$. Therefore, potentially, in the last $m-t$ coordinates of $G$ we can have terms of the form $x_{i}$, $y_{i}$, $x_{i}x_{j}$, $x_{i}y_{j}$, $y_{i}y_{j}$, $x_{i}y_{j}y_{k}$, $y_{i}y_{j}y_{k}$, $y_{i}y_{j}y_{k}y_{l}$. Notice that, even if we replace $U$ with a dense quadratic map, the possible terms in $G$ have the same form. To sum up, even if the degree of $G$ is up to 4, the terms of degree 4 involve only variables in $y$, while terms of degree 3 involve only variables in $y$ or 2 variables in $y$ and 1 variable in $x$.
The public map $G_{pub}=A_{1}\circ G\circ A_{2}$ consists of dense polynomials of degree up to 4, since the random map $A_{2}$ will remove the above-mentioned restrictions. However, the affine transformation keeps invariant the amount of components of a fixed degree, so to analyse the possible degrees of the components of $G_{pub}$, we can directly consider the map $G(x,y)=(x-{\mathfrak{q}}(y),U(x-{\mathfrak{q}}(y),y))$. So, $G_{pub}$ has at least $q^{t}-1$ quadratic components, since the components $\lambda\cdot G$ with $\lambda=(v,0_{m-t})$ with nonzero $v\in{\mathbb{F}}_{q}^{t}$ are quadratic. Moreover, we note that in all the systems constructed during the preparation of this paper, $U(x-{\mathfrak{q}}(y),y)$ does not contain quadratic components. This is expected since the polynomials in $U$ and ${\mathfrak{q}}$ are quadratic polynomials with random coefficients. In §6.2, we will show how this can be leveraged to propose a reduction in the key sizes.
@@ 5.2 Dimensions of the keys
In the following, we provide an analysis of the sizes of the public and private keys. Furthermore, in Table 1 we will present a collection of concrete values of the key sizes for a selection of parameters.
@@
Recall that a polynomial of degree $r$ in $n$ variables with coefficients over ${\mathbb{F}}_{q}$ has $M_{n}(r)={n+r\choose r}$ terms. Notice that when $q\leq r$, we can use the field equations $x_{i}^{q}-x_{i}$ to lower the degrees of the polynomials over ${\mathbb{F}}_{q}$ and store fewer coefficients. $G_{pub}$ is a system of $m$ equations of degree at most 4 in $n$ variables, where each equation has $M_{n}(4)={n+4\choose 4}$ possible terms. Hence, for $G_{pub}$ we need to specify $m\cdot M_{n}(4)$ coefficients over ${\mathbb{F}}_{q}$.
@@
  • $A_{1}$ and $A_{2}$ are affine bijections over ${\mathbb{F}}_{q}^{m}$ and ${\mathbb{F}}_{q}^{n}$ respectively: They correspond to invertible $m\times m$ and $n\times n$ matrices over ${\mathbb{F}}_{q}$ plus a constant vector. So this means $m^{2}+m+n^{2}+n$ coefficients over ${\mathbb{F}}_{q}$.
@@ 5.3 Computational cost of the procedure
In this subsection, we present a tentative analysis of the computational cost of the proposed multivariate scheme. Notice that the application of this procedure reduces to the evaluation of polynomials of degree at most four and computing the solution of linear systems. Considering the direct evaluation of a polynomial, we have the following estimates.
@@
This is clearly an upper bound, since more efficient techniques might be used, see for example [BES13]. We indicate with $\mathfrak{M}(r,n)$ the number of multiplications needed to solve a linear system of $r$ equations in $n$ variables. Recall that $\mathfrak{M}(r,n)$ is at most of the order $rn\cdot\min(r,n)$, see for example [BV18, Appendix B].
@@
Consider a signature scheme based on the Pesto scheme as in Definition 4.2. Set $\mathfrak{m}_{i}(r)$ and $\mathfrak{M}(r,k)$ as defined before. Excluding the computation of the hash value, the computational cost, in terms of multiplications over ${\mathbb{F}}_{q}$, to verify the validity of a signature is $m\cdot\mathfrak{m}_{4}(n)$, whereas the computational cost to produce a valid signature is $m\cdot\mathfrak{m}_{1}(m)+t\cdot\mathfrak{m}_{2}(n-t)+n\cdot\mathfrak{m}_{1}(n)+(m-t)\left(\mathfrak{m}_{2}(t)+\mathfrak{m}_{2}(s)\right)+\mathfrak{M}(m-t,n-t-s)$.
@@
We are left with analysing the cost for the third step, where we have to find a preimage of $U(w_{T}^{\prime},y)=w_{U}^{\prime}$. Hence, we need to set $y_{1},\ldots,y_{s}$ to random values, evaluate those variables and then solve a linear system. The cost of evaluating $s$ variables corresponds to $(m-t)\cdot\mathfrak{m}_{2}(s)=(m-t)s(s+2)$ multiplications. The cost of solving the final system of $m-t$ equations in $n-t-s$ variables is $\mathfrak{M}(m-t,n-t-s)$ multiplications.
@@ 6 Security Analysis
In this section, we provide some considerations on the security of the scheme Pesto of Definition 4.2. Among them, we consider attacks that have been exploited for the MI cryptosystem (and its generalizations) and we analyse under which conditions it would be possible to extend these also to our system.
@@ 6.1 The importance of $A_{2}$
We start this analysis by presenting a first observation on the security of the scheme, which stresses the importance of the affine bijection $A_{2}$. Indeed, if an attacker is able to recover $A_{2}$, then they can operate as follows. First, compute $\bar{G}=G_{pub}\circ A_{2}^{-1}$ and isolate its quadratic components: $t$ of them are of the form $x_{i}-{\mathfrak{q}}_{i}(y)$, for $i=1,\ldots,t$. From this, it is rather easy to solve the system. Suppose a ciphertext $c$ is given, and the attacker wants to find a message $m$ such that $G_{pub}(m)=c$. Write $G_{pub}(m)=\bar{G}\circ A_{2}(m)=c$. From the isolated quadratic equations, recover $x={\mathfrak{q}}(y)+c_{T}$. By substituting this into the remaining equations of $\bar{G}$, solve the system, which has at most $s$ variables appearing in quadratic terms. Notice that this last step is equivalent to what a legitimate user has to do. Once the solution $(x,y)$ of this system is found, then $m=A_{2}^{-1}(x,y)$.
In order to perform this attack, the attacker has to completely recover $A_{2}$ but only partially $A_{1}$ (it is enough to isolate the quadratic components of the form $x_{i}-{\mathfrak{q}}_{i}(y)$). The cost to isolate these quadratic components will be treated in the following section.
6.2 Isolating the quadratic components: the shape of $A_{1}$
In this section, we explain a strategy to isolate, within $G_{pub}$, the components that correspond to a (linear combination of) the first $t$ coordinates of $G$. This demonstrates that, in many situations, the public system can be partitioned into two sets of equations and suggests that the external affine transformation $A_{1}$ can be chosen in a specific form, thereby reducing the size of the secret key.
We try to estimate the cost of isolating the quadratic components of $G_{pub}$. Let $r\geq t$ be such that the number of quadratic components of $G_{pub}$ is $q^{r}-1$. Notice that when $r=t$, then $U(x-{\mathfrak{q}}(y),y)$ does not have quadratic components. As mentioned in §5.1, this situation occurred in all the experiments we conducted. We consider the list $\Delta$ of all possible terms of degree 3 and degree 4 in $n$ variables appearing in $G_{pub}$. We have
@@ 6.2 Isolating the quadratic components: the shape of $A_{1}$
We construct the matrix $A$ with $n$ rows and $|\Delta|$ columns, with at position $(i,j)$ the coefficients of the $j$-th term of $\Delta$ in the $i$-th equation of $G_{pub}$. So, to recover the quadratic equations of $G_{pub}$ it is enough to solve the linear system $x^{T}A=A^{T}x=0_{|\Delta|}$. The set of solutions corresponds to the set of quadratic components of $G_{pub}$. The size of this set is $q^{r}$. Recall that the cost of solving a linear system with $n$ equations in $|\Delta|$ variables $\mathfrak{M}(n,|\Delta|)$ is of the order $O(n^{6})$. Note, however, that this approach allows us to recover the space of quadratic components of $G\circ A_{2}$, but not the exact quadratic coordinates.
Remark 6.1.
The previous observations suggest that a dense affine transformation $A_{1}$, which completely mixes the first $t$ equations of $G$ with the remaining $m-t$ equations, may not provide additional security, since the first $t$ equations can be recovered at a cost of $O(n^{6})$ by an attacker. Therefore, we suggest considering the matrix corresponding to the linear part of $A_{1}$ in the form $\begin{pmatrix}*&0\\ *&*\end{pmatrix}$, with a zero block of size $t\times(m-t)$. This way, we do not fully mix all the equations and we reduce the secret key and public key sizes. With this observation, the secret key size of Pesto from Proposition 5.2 becomes
$m^{2}-t(m-t)+m+n^{2}+n+t{n-t+2\choose 2}+(m-t){t+s+2\choose 2}+(m-t)(n-t-s)(t+s+1)$
coefficients over ${\mathbb{F}}_{q}$ to store; the public key size of Pesto from Proposition 5.2 becomes
$(m-t)\cdot{n+4\choose 4}+t\cdot{n+2\choose 2}$
coefficients over ${\mathbb{F}}_{q}$ to store.
In Table 1, we present a comparison of the sizes of secret and public keys between the original version with a generic affine transformation $A_{1}$ as in Proposition 5.2 and this modified version. The choice of parameters will be explained in §6.7.
$n$ $m$ $t$ $s$ amount for $pk$ amount for $sk$ amount for $pk_{r}$ amount for $sk_{r}$
27 25 10 2 $786$\cdot 10^{3}$ $7406$ $476$\cdot 10^{3}$ $7256$
40 38 14 2 $515$\cdot 10^{4}$ 21 8782187821\,87821 878 $327$\cdot 10^{4}$ 21 5422154221\,54221 542
57 55 20 2 $287$\cdot 10^{5}$ 59 0415904159\,04159 041 $182$\cdot 10^{5}$ 58 3415834158\,34158 341
Table 1: Amount of coefficients of $\mathbb{F}_{q}$ to store, where $pk$ and $sk$ refer to the original system, and $pk_{r}$ and $sk_{r}$ refer to the reduced system with $A_{1}$ with a zero block.
6.3 Using linear structures to analyse $A_{2}$
In this section, we present an analysis using linear structures (see Definition 6.2 below for the precise definition) of some components of $G_{pub}$. We outline the core idea and explain how, under suitable assumptions, this can be used to recover partial information on the affine transformation $A_{2}$. However, at this stage, this analysis does not appear sufficient to fully determine $A_{2}$ and, consequently, does not seem to compromise the scheme.
@@ 6.3 Using linear structures to analyse $A_{2}$
    $G_{pub}$ has (at least) $q^{t}-1$ quadratic components with (at least) $q^{t}$ linear structures in common. Specifically, they are the linear combinations of the first $t$ coordinates of $G\circ A_{2}$. Indeed, for quadratic functions of the form $x_{i}+{\mathfrak{q}}_{i}(y)$, and their linear combinations, there are at least $q^{t}$ linear structures common to all these functions.
@@
    If we are able to identify this subspace, we know that the linear part of $A_{2}$ maps $V$ into ${\mathbb{F}}_{q}^{t}$. So, we can recover some partial information on this linear part. Specifically, we know that in $V$ there are $t$ linearly independent elements which form the first $t$ columns of the inverse of the linear part of $A_{2}$.
@@
The multiset of the number of linear structures of the components of $G$ and of $G_{pub}$ is the same. Moreover, $a\in{\mathbb{F}}_{q}^{n}$ is a linear structure of $\lambda\cdot G_{pub}$ if and only if $L_{2}(a)$ is a linear structure of ${L_{1}^{*}(\lambda)}\cdot G$ where $L_{1}$ and $L_{2}$ are the linear parts of the affine bijections $A_{1}$ and $A_{2}$, and $L_{1}^{*}$ is the adjoint operator of $L_{1}$, that is $x\cdot L_{1}(y)=L_{1}^{*}(x)\cdot y$ for every $x,y$.
Proof 6.6.
@@
The second statement can be deduced from the following. From the definition of $G_{pub}$ and since $A_{1}$ is affine we have $D_{a}(\lambda\cdot G_{pub})(x)=\lambda\cdot G_{pub}(x+a)-\lambda\cdot G_{pub}(x)=\lambda\cdot A_{1}\circ G\circ A_{2}(x+a)-\lambda\cdot A_{1}\circ G\circ A_{2}(x)=\lambda\cdot L_{1}\circ G\circ A_{2}(x+a)-\lambda\cdot L_{1}\circ G\circ A_{2}(x)$. Then $D_{a}(\lambda\cdot G_{pub})(x)=\lambda\cdot L_{1}(G(A_{2}(x)+L_{2}(a))-G(A_{2}(x)))=\lambda\cdot L_{1}(D_{L_{2}(a)}G(A_{2}(x)))=L_{1}^{*}(\lambda)\cdot D_{L_{2}(a)}G(A_{2}(x))$.
@@
The desired components of $G_{pub}$ are quadratic; therefore by isolating all the quadratic components of the public function, we are isolating the desired components plus (eventually) the quadratic components of $U(x-{\mathfrak{q}}(y),y)$. Recall that in all the computational experiments conducted, $U(x-{\mathfrak{q}}(y),y)$ did not have components of degree 2.